# On the ring of cooperations for Real Hermitian K-theory

### v1

@@ 7.1. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
The group $\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ is zero in coweight $cw\equiv 1\,(4)$.
@@
In coweight $cw\equiv 1\,(4)$, the $E_{\infty}$-page of the aAHSS$(B_{0}^{\mathbb{R}}(1))$ is isomorphic to the a direct sum of coweight $cw\equiv 1\,(4)$ in Atiyah–Hirzebruch filtration 0, coweight $cw\equiv 0\,(4)$ in Atiyah–Hirzebruch filtration 2, and coweight $cw\equiv 3\,(4)$ in Atiyah–Hirzebruch filtration 3 pieces. The differentials:
@@
ensure that Atiyah–Hirzebruch filtration 2 is trivial in coweight $cw\equiv 0\,(4)$. Atiyah–Hirzebruch filtration 3 is trivial in coweight $cw\equiv 3\,(4)$ since $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is trivial in this coweight. ∎
We organize the $E_{\infty}$-page by coweight modulo 4 in the same way that we depicted $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$. As indicated by Lemma 7.2, the different Atiyah–Hirzebruch filtration pieces contribute different pieces to each coweight of the $E_{\infty}$-page, suitably shifted by the degree of the cell in the filtration it is attached to. Since $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is concentrated in coweights $cw\equiv 0,1,\text{ and }2\,(4)$, for each Atiyah–Hirzebruch filtration piece there is a coweight on the $E_{\infty}$-page which it does not contribute to. To make this clear, we describe each coweight piece of the $E_{\infty}$-page in terms of contributions from Atiyah–Hirzebruch filtration in Table 3.
$E_{\infty}$ coweight Filtration 0 Filtration 2 Filtration 3
@@
With this in mind, we are prepared to give charts for the $E_{\infty}$-page. Fig. 18 depicts the $E_{\infty}$-page in coweight $cw\equiv 0\,(4)$. Fig. 19 depicts the $E_{\infty}$-page in coweight $cw\equiv 2\,(4)$. Fig. 20 depicts the $E_{\infty}$-page in coweight $cw\equiv 3\,(4)$. We use the same notation as previously used in this section, with some additions. A $\square$ represents $\mathbb{F}_{2}[\tau^{4}][\rho]$, where in contrast to a black $\blacksquare$, the $\rho$-tower does not support $h_{1}$-multiplication. Red lines indicate hidden extensions. A diagonal black $\nearrow$ indicates infinite $h_{1}$-multiplication on a $\rho$-tower which is compatible with an infinite $\rho$-tower off of a class which is $h_{1}$-divisible. For example, in coweight $cw\equiv 2\,(4)$, there is a class $x$ in degree $(3,1)$ which supports an infinite $h_{1}$-tower, and there is a class $y$ in degree $(5,2)$ which is part of an infinite $h_{1}$-tower and supports an infinite $\rho$-tower. Hidden in our charts, and as indicated by $\nearrow$, is the relation $h_{1}x=\rho y.$
@@
The big flag $\mathcal{F}_{4,1,2}$ is isomorphic to the coweight $cw\equiv 2\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 19).
We note that most classes on the $E_{\infty}$-page are $v_{1}$-periodic. In fact, the only classes which are $v_{1}$-torsion are those in the $\rho$-tower in Adams filtration 0 of the coweight $cw\equiv 0\,(4)$ piece which are divisible by $\rho^{3}$.
@@
There is an isomorphism
$\textup{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(B^{\mathbb{R}}_{0}(1))\cong\pi_{**}^{\mathbb{R}}(\textup{ksp}).$
@@
7.2. $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ and recursive submodules
@@
To illustrate the numerous $v_{1}$-torsion classes which arise in the process of computing $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, which are ultimately to be ignored (see Remark 4.3), we explicitly compute $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$. We form an algebraic Atiyah–Hirzebruch spectral sequence by taking the cellular filtration on $B_{0}^{\mathbb{R}}(1)$ from before and applying the functor $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)\otimes-\right)$. This has signature
$E_{1}=\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))\otimes\mathbb{M}_{2}^{\mathbb{R}}\{1,\bar{\xi_{1}},\bar{\tau_{1}}\}\implies\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2}).$
@@
The differentials are determined in an analogous manner as before, and hidden extensions can be realized by the same methods. Similar to the previous case, the behavior of the differentials ensures that the $E_{\infty}$-page is trivial in coweight $cw\equiv 1\,(4)$. We present the $E_{\infty}$ page, modulo $v_{1}$-torsion, below. Fig. 21 depicts the $E_{\infty}$-page in coweight $cw\equiv 0\,(4)$. Fig. 22 depicts the $E_{\infty}$-page in coweight $cw\equiv 1\,(4)$. Fig. 23 depicts the $E_{\infty}$-page in coweight $cw\equiv 2\,(4)$. Note that we do not present hidden extensions in red in these charts.
Figure 21. The $E_{\infty}$-page of the aAHSS$\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ in $cw\equiv 0\,(4)$, modulo $v_{1}$-torsion.
Figure 22. The $E_{\infty}$-page of the aAHSS$\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ in $cw\equiv 1\,(4)$, modulo $v_{1}$-torsion.
Figure 23. The $E_{\infty}$-page of the aAHSS$\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ in $cw\equiv 2\,(4)$, modulo $v_{1}$-torsion.
@@
The big flag $\mathcal{F}_{8,2,4}$ is isomorphic to the summand of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ concentrated in stems $s\geq 2$ (see Fig. 21).
Before giving a general formula for $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, we make a few simple observations. The group $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, is only nonzero in coweights $cw\equiv 0,1,2\,(4)$. After running the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)$), we see that there is a submodule
$M_{1}:=\Sigma^{4,2}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle 1\rangle\subseteq\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)),$
using the notation from Section 5. Note that if $x$ has coweight $cw\equiv t\,(4),$ then $\Sigma^{4,2}x\langle 1\rangle$ has coweight $cw\equiv t+2\,(4)$. In particular, $M_{1}$ is concentrated in coweights $cw\equiv 0,2,3\,(4)$. By observation, we see that in fact all classes in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ are concentrated in these coweights. Similarly, after running the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ we see that there is a submodule
$M_{2}:=\Sigma^{8,4}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle 2\rangle\subseteq\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right).$
It follows that $M_{2}$ is concentrated in coweights $cw\equiv 0,1,2\,(4)$, and by observation we see that in fact all classes in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ are concentrated in these coweights. More generally, from the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ we see that the group $\text{Ext}_{\mathcal{A}_{\mathbb{R}}^{\vee}}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$ contains a submodule
$M_{i}=\Sigma^{4i,2i}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle i\rangle\subseteq\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right),$
and that, modulo $v_{1}$-torsion, $\text{Ext}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ is concentrated in the same coweights as $M_{n}$ is. More precisely, we have the following result.
@@
[…]
We also have the following submodule of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i}).$
@@
[…]
$\Sigma^{4i,2i}\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)\langle i\rangle\subseteq\textup{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right).$
@@
We have shown this to be true in the case of $i=1$ in Corollary 7.4. The differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ on the submodule $M_{1}$ are isomorphic to the differentials in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$, hence the submodule $M_{2}$ of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ extends to give a submodule
$\Sigma^{8,4}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)\right)\langle 2\rangle\subseteq\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right).$
For arbitrary $i$, the submodule $M_{i-1}\subseteq\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i-1}\right)$ extends to the submodule
$\Sigma^{4(i-1),2(i-1)}\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)\langle i-1\rangle$
for each $11\leq j\leq i-$ 1. The differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ are isomorphic to the differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)$ when restricted to this submodule, hence the result of the spectral sequence gives the desired submodules in the target. ∎

### v2

@@ 7.1. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$
The group $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ is zero in coweight $cw\equiv 1\,(4)$.
@@
[…]
@@
ensure that Atiyah–Hirzebruch filtration 2 is trivial in coweight $cw\equiv 0\,(4)$. Atiyah–Hirzebruch filtration 3 is trivial in coweight $cw\equiv 3\,(4)$ since $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is trivial in this coweight. ∎
We organize the $\textup{E}_{\infty}$-page by coweight modulo 4 in the same way that we depicted $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$. As indicated by Lemma 7.2, the different Atiyah–Hirzebruch filtration pieces contribute different pieces to each coweight of the $\textup{E}_{\infty}$-page, suitably shifted by the degree of the cell in the filtration it is attached to. Since $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$ is concentrated in coweights $cw\equiv 0,1,\text{ and }2\,(4)$, for each Atiyah–Hirzebruch filtration piece there is a coweight on the $\textup{E}_{\infty}$-page which it does not contribute to. To make this clear, we describe each coweight piece of the $\textup{E}_{\infty}$-page in terms of contributions from Atiyah–Hirzebruch filtration in Table 3.
$\textup{E}_{\infty}$ coweight Filtration 0 Filtration 2 Filtration 3
@@
With this in mind, we are prepared to give charts for the $\textup{E}_{\infty}$-page. Figure 7.9 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 0\,(4)$. Figure 7.10 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 2\,(4)$. Figure 7.11 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 3\,(4)$. We use the same notation as previously used in this section, with some additions. A $\square$ represents $\mathbb{F}_{2}[\tau^{4}][\rho]$, where in contrast to a black $\blacksquare$, the $\rho$-tower does not support $h_{1}$-multiplication. For example, in coweight $cw\equiv 0\,(2)$, there is a class $z$ in degree (1,0)(-1,0) which is a $\rho$-multiple of the class in degree $(0,0)$. As the symbol at $(0,0)$ is a $\square$, the class $z$ is $h_{1}$-torsion. Red lines indicate hidden extensions. A diagonal black $\nearrow$ indicates infinite $h_{1}$-multiplication on a $\rho$-tower which is compatible with an infinite $\rho$-tower off of a class which is $h_{1}$-divisible. For example, in coweight $cw\equiv 2\,(4)$, is a relation
$h_{1}x=\rho y,$
where $x$ is the class in degree $(3,1)$ which supports an infinite $h_{1}$-tower, and $y$ is the class in degree $(5,2)$ which supports an infinite $\rho$-tower. This relation occurs in some $\mathbb{F}_{2}$ which is not present in the current chart: the current $\mathbb{F}_{2}$ in degree $(4,2)$ is a part of an $h_{0}$-tower, and thus is not $h_{1}$-divisible. We suppress notation and indicate this relation by $\nearrow$ rather than drawing an extra $\bullet$.
@@
The big flag $\mathcal{F}_{4,1,2}$ is isomorphic to the coweight $cw\equiv 2\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ (see Fig. 7.10).
We note that most classes on the $\textup{E}_{\infty}$-page are $v_{1}$-periodic. In fact, the only classes which are $v_{1}$-torsion are those in the $\rho$-tower in Adams filtration 0 of the coweight $cw\equiv 0\,(4)$ piece which are divisible by $\rho^{3}$.
@@
The $\textup{{mASS}}^{\mathbb{R}}(\textup{ksp})$ has signature
$\textup{E}_{2}^{s,f,w}=\textup{Ext}^{s,f,w}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}(B_{0}^{\mathbb{R}}(1))\implies\pi^{\mathbb{R}}_{s,w}(\textup{ksp})$
and collapses on the $\textup{E}_{2}$-page.
@@
7.2. $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ and recursive submodules
@@
To illustrate the numerous $v_{1}$-torsion classes which arise in the process of computing $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, which are ultimately to be ignored (see Remark 4.2), we explicitly compute $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$. We form an algebraic Atiyah–Hirzebruch spectral sequence by applying the functor $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)\otimes-\right)$ to (4.3). This has signature
$\textup{E}_{1}=\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}\implies\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2}).$
The differentials are again induced by the algebraic cell structure of $B_{0}^{\mathbb{R}}(1)$. Thus, the $d_{1}$-differential is only nonzero on Atiyah–Hirzebruch filtration 3, where
$d_{1}(\alpha[3])=h_{0}\alpha[2],$
the $d_{2}$-differential is only nonzero on Atiyah–Hirzebruch filtration 2, where
$d_{2}(\alpha[2])=h_{1}\alpha[0],$
and the $d_{3}$-differential is only nonzero on Atiyah–Hirzebruch filtration 3, where
$d_{3}(\alpha[3])=\langle\alpha,h_{0},h_{1}\rangle[0].$
Since the $d_{1}$- and $d_{2}$-differentials are given by a product with an element of $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(\mathbb{M}_{2}^{\mathbb{R}})$ of coweight 0, they preserve coweight between Atiyah-Hirzebruch filtrations. Similar to the previous case, the $d_{3}$-differential increases coweight by 1.
In fact, most of the work in running the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ follows immediately from our previous computation of the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ in Section 7.1. By observation, we see that there is a submodule
$M_{1}:=\Sigma^{4,2}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle 1\rangle\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)),$
using the notation from Section 5. This submodule is very large in $\text{Ext}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}^{***}(B_{0}^{\mathbb{R}}(1))$: in coweight $cw\equiv 0\,(4)$, $M_{1}$ consists of everything except for the $h_{1}$-torsion $(\rho,h_{0})$-tower in stem 0; in coweight $cw\equiv 2\,(4)$, $M_{1}$ consists of everything except for the $(\rho,h_{0})$-tower in stem 0; and in coweight $cw\equiv 3\,(4)$, it consists of every nonzero element.
As a result, there is a submodule of the $\textup{E}_{1}$-page of the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ given by
(7.2) $M_{1}\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}.$
The differentials on this submodule are identical to the differentials in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$. Thus, one can depict the differentials in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$ on the summand (7.2) by shifting the charts given in Figure 7.1 through Figure 7.6 to the appropriate tridegree and coweight. The hidden extensions on this summand are the exact same as those in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$.
@@
The differentials on the classes not in the $M_{1}\otimes\mathbb{M}_{2}^{\mathbb{R}}\{[1],[\overline{\xi}_{1}],[\overline{\tau}_{1}]\}$ are straightforward to compute by the formulas given, as are hidden extensions. Similar to the previous case, the behavior of the differentials ensures that the $\textup{E}_{\infty}$-page is trivial in coweight $cw\equiv 1\,(4)$. We present the $\textup{E}_{\infty}$ page, modulo $v_{1}$-torsion, below. Figure 7.12 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 0\,(4)$. Figure 7.13 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 1\,(4)$. Figure 7.14 depicts the $\textup{E}_{\infty}$-page in coweight $cw\equiv 2\,(4)$. Note that we do not present hidden extensions in red in these charts.
[…]
@@
The big flag $\mathcal{F}_{8,2,4}$ is isomorphic to the summand of the coweight $cw\equiv 0\,(4)$ piece of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes 2})$ concentrated in stems $s\geq 2$ (see Figure 7.12).
Before giving a general formula for $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$, we make a few simple observations. The group $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})$, is only nonzero in coweights $cw\equiv 0,1,2\,(4)$. After running the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1)$), we saw that there was a submodule
$M_{1}:=\Sigma^{4,2}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle 1\rangle\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)),$
Note that if $x$ has coweight $cw\equiv t\,(4),$ then $\Sigma^{4,2}x\langle 1\rangle$ has coweight $cw\equiv t+2\,(4)$. In particular, $M_{1}$ is concentrated in coweights $cw\equiv 0,2,3\,(4)$. By observation, we see that in fact all classes in $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1))$ are concentrated in these coweights. Similarly, after running the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ we see that there is a submodule
$M_{2}:=\Sigma^{8,4}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle 2\rangle\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right).$
It follows that $M_{2}$ is concentrated in coweights $cw\equiv 0,1,2\,(4)$, and by observation we see that in fact all classes in $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ are concentrated in these coweights. More generally, from the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ we see that the group $\text{Ext}^{***}_{\mathcal{A}_{\mathbb{R}}^{\vee}}(B_{0}^{\mathbb{R}}(1)^{\otimes i})$ contains a submodule
$M_{i}=\Sigma^{4i,2i}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(\mathbb{M}_{2}^{\mathbb{R}})\langle i\rangle\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right),$
and that, modulo $v_{1}$-torsion, $\text{Ext}^{***}_{\mathcal{A}(1)_{\mathbb{R}}^{\vee}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ is concentrated in the same coweights as $M_{i}$ is. More precisely, we have the following result.
@@
The group $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)/v_{1}\textup{-torsion}$ is trivial in coweight $cw\equiv 1\,(4)$ for $i\equiv 1\,(2)$ and in coweight $cw\equiv 3\,(4)$ for $i\equiv 0\,(2)$.
We also have the following submodule of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}(B_{0}^{\mathbb{R}}(1)^{\otimes i}).$
@@
Let $i\geq 1$. For all $1\leq j\leq i$, the group $\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)/v_{1}\textup{-torsion}$ contains as a submodule:
$\Sigma^{4i,2i}\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)\langle i\rangle\subseteq\textup{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right).$
@@
We have shown this to be true in the case of $i=1$ in Corollary 7.3. The differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ on the submodule $M_{1}$ are isomorphic to the differentials in the $\textbf{aAHSS}(B_{0}^{\mathbb{R}}(1))$, hence the submodule $M_{2}$ of $\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right)$ extends to give a submodule
$\Sigma^{8,4}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)\right)\langle 2\rangle\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes 2}\right).$
For arbitrary $i$, the submodule $M_{i-1}\subseteq\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i-1}\right)$ extends to the submodule
$\Sigma^{4(i-1),2(i-1)}\text{Ext}^{***}_{\mathcal{A}(1)^{\vee}_{\mathbb{R}}}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)\langle i-1\rangle$
for each $1\leq j\leq i-1$. The differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes i}\right)$ are isomorphic to the differentials in the $\textbf{aAHSS}\left(B_{0}^{\mathbb{R}}(1)^{\otimes j}\right)$ when restricted to this submodule, hence the result of the spectral sequence gives the desired submodules in the target. ∎